### v1

# Neuromodulation and homeostasis: complementary mechanisms for robust neural function

ABSTRACT
Neurons depend on two interdependent mechanisms—homeostasis and neuromodulation—to maintain robust and adaptable functionality. Homeostasis stabilizes neuronal activity by adjusting ionic conductances, whereas neuromodulation dynamically modifies ionic properties in response to external signals. Combining these mechanisms in conductance-based models often produces unreliable outcomes, particularly when sharp neuromodulation interferes with homeostatic tuning. This study explores how a biologically inspired neuromodulation controller can harmonize with homeostasis to ensure reliable neuronal function. Using computational models of stomatogastric ganglion and dopaminergic neurons, we demonstrate that controlled neuromodulation preserves neuronal firing patterns while maintaining intracellular calcium levels. Unlike sharp neuromodulation, the neuromodulation controller integrates activity-dependent feedback through mechanisms mimicking G-protein-coupled receptor cascades. The interaction between these controllers critically depends on the existence of an intersection in conductance space, representing a balance between target calcium levels and neuromodulated firing patterns. Maximizing neuronal degeneracy enhances the likelihood of such intersections, enabling robust modulation and compensation for channel blockades. We further show that this controller pairing extends to network-level activity, reliably modulating central pattern generators in crustaceans. These findings suggest that targeting neuromodulation pathways—rather than ion channels directly—may offer safer pharmacological strategies to manage neuronal dysfunctions. This study highlights the complementary roles of homeostasis and neuromodulation, proposing a unified control framework for maintaining robust and adaptive neural activity under physiological and pathological conditions.
SIGNIFICANCE STATEMENT
Neurons must maintain stable activity while adapting to changing demands, relying on two mechanisms: homeostasis, which stabilizes activity, and neuromodulation, which adjusts behavior to external signals. These mechanisms often interact, but their improper coordination can lead to dysfunction. This study uses computational models to show how controlled neuromodulation - mimicking biological feedback - can be harmonized with homeostasis to ensure robust neuronal function. By improving neuronal degeneracy, the system compensates for channel blockades and maintains critical activity patterns for both single neurons and networks. These insights not only deepen our understanding of neural robustness, but also suggest safer pharmacological strategies targeting neuromodulatory pathways, offering potential breakthroughs in treating neurological disorders without disrupting essential cellular mechanisms.
INTRODUCTION
Brain activity is continuously shaped by neuromodulators and neuropeptides, including dopamine, serotonin, and histamine (Bargmann and Marder,, 2013; Marder et al.,, 2014). Neuromodulators dynamically influence single-neuron activity, input-output properties, and synaptic connection strength and dynamics, enabling neuronal networks to adapt to changing needs, contexts, and environments (Marder and Calabrese,, 1996; Marder and Bucher,, 2001; McCormick et al.,, 2020). To achieve this modulation, neuromodulators dynamically reshape the density, dynamics, and kinetics of many single cell transmembrane proteins, making whole brain functional signaling strongly dependent on the robustness and reliability of neuromodulation mechanisms at the molecular and cellular levels (Nadim and Bucher,, 2014; Hooper and Marder,, 1984). While experimental studies underscore the ubiquity of neuromodulation across nervous systems, the fundamental principles underlying this mechanism remain incompletely understood. As such, studying neuromodulation is a very arduous task, both experimentally and computationally, which explains our limited understanding of its underlying mechanisms.
Neuromodulation coexists with cellular homeostasis, a mechanism that gradually adjusts neuronal membrane properties to maintain a target activity level (Desai,, 2003; Marder and Goaillard,, 2006; O’Leary et al.,, 2013). Neuronal cells are dependent on homeostatic mechanisms to maintain optimal functionality throughout the extended useful life of mammals (Marder,, 2011). Despite continuous turnover of transmembrane proteins, such as ion channels and receptors, occurring on varying temporal scales ranging from hours to weeks (O’Leary and Wyllie,, 2011), the robustness and excitability of neurons must remain undisturbed (Marder and Prinz,, 2002). By monitoring only intracellular calcium levels, the neuron orchestrates the regulation of all its ion channel conductances through homeostatic self-tuning rules (O’Leary et al.,, 2010; Pratt and Aizenman,, 2007). This raises the question of how two different mechanisms acting on the same targets (ion channels), each potentially having divergent goals, can co-exist and complement each other harmoniously (Marder et al.,, 2014; Marder and Prinz,, 2002).
We propose to study this interaction between neuromodulation and homeostasis computationally, using conductance-based models. Conductance-based models provide an electrical representation of the neuronal membrane, where the membrane is modeled as an RC circuit, while the ion channels are modeled as non-linear conductances (Izhikevich,, 2007). These models are the closest equivalent of biological neurons, thanks to their ability to model any ion channel from experimental data following a methodology first proposed by Hodgkin and Huxley (Hodgkin and Huxley,, 1952). There exists a myriad of different conductance-based models that model various types of neuronal cells (Liu et al.,, 1998; Qian et al.,, 2014; Plant,, 1981; Wechselberger et al.,, 2006; Drion et al.,, 2011; Katori et al.,, 2010; Megwa et al.,, 2023).
On the one hand, a homeostatic controller has been developed by O’Leary et al., (2014) and can be applied on top of any conductance-based model. In this controller, the neuron can adjust all ionic conductances up or down depending on its average intracellular calcium level. If the calcium level is lower than a reference value, ionic conductances are increased following homeostatic tuning rules and vice versa. Such rules force the tuning of ionic conductances along a single axis, that is, by preserving conductance ratios or correlations. On the other hand, Fyon et al., (2023) has developed a reliable neuromodulation controller and can also be applied to any conductance-based model. This controller can robustly tune the external activity of neurons, which means that neurons with various sets of ion channel properties that lead to similar firing activities (a property called degeneracy (Drion et al., 2015b, )) can be neuromodulated using the same control action. In this study, we first examine the interaction between homeostasis and sharp neuromodulation, emphasizing that improperly combining the homeostatic controller with neuromodulation may result in unreliable outcomes. Subsequently, we expand our analysis to explore the interaction between the homeostatic controller and a neuromodulation controller, both at the single-cell level and in rhythmic networks. This approach enables the development of a combined controller that integrates the strengths of both individual controllers, facilitating robust modulation and preservation of neuronal function.
RESULTS
Current model of homeostasis can lead to pathological excitability in the presence of sharp neuromodulation
This study employs a stomatogastric ganglion (STG) neuron conductance model (Liu et al.,, 1998) to explore the interaction between homeostasis and neuromodulation. These neurons contribute to rhythmic circuits that generate signals triggering stomach muscle contractions in crustaceans. This model was chosen for its ability to produce diverse firing patterns—ranging from tonic spiking to bursting, including single-spike bursting—and its rich combination of ion channels enabling degeneracy. The model includes classical fast sodium and delayed rectified potassium channels, A-type and calcium-activated potassium channels, T-type and slow calcium channels as well as hyperpolarization-activated H-channels. We first study the interaction between state-of-the-art model homeostasis (O’Leary et al.,, 2014) and sharp neuromodulation (Fyon et al.,, 2024). We define sharp neuromodulation as a direct change in a subset of ion channel maximal conductances, which results in functional modulation. In the context of this work, we will focus on a robust transition from tonic spiking to bursting in degenerate neurons, i.e. neurons exhibiting similar firing activities from different underlying channel conductances.
First, we sharply neuromodulated a set of degenerate neurons from tonic spiking to bursting by directly modifying the maximal conductances of A-type potassium $\bar{g}_{A}$ and slow calcium $\bar{g}_{CaS}$, then allowed homeostatic compensation to take its course. Three example traces with a time evolution of maximal conductances are provided in Fig. 1, left, whereas Fig. 1, right show trajectories in the $\bar{g}_{\mathrm{CaS}}$-$\bar{g}_{\mathrm{A}}$ plane. As can be seen in the figure, although the initial tonic firing and the sharply neuromodulated bursting activities are very similar in all three neurons, homeostatic compensation has very variable and unreliable effects. Depending on the initial conductance state, neuromodulation shifts the neuron to a new state where homeostatic compensation can lead to pathological (Fig. 1A), functionally preserved (Fig. 1B), or unphysiological (Fig. 1C) behaviors. This interesting result highlights an important property of neuronal degeneracy: although distinct channel subsets can produce similar neuronal behaviors, neurons may respond differently to perturbations, such as sharp neuromodulation.
To better understand the reasons behind this unreliability, both mechanisms must be discussed in detail. In the homeostatic model O’Leary et al., (2014), the cell senses its own intracellular calcium level and averages it over time. If the mean intracellular calcium falls below a predefined target value, RNA transcription and ion channel protein translation are triggered to increase all ionic conductance values, and vice versa. Specifically, increasing calcium ion channel conductances leads to more calcium ions entering the cell. This implies that homeostasis induces changes in all conductance values (Fig. 1 left), which leads to a linear motion in the conductance plane (Fig. 1 right). This phenomenon happens on a relatively slow timescale, which can last up to days (O’Leary et al.,, 2010). Notably, the RNA transcription time constants, denoted by $\tau_{m}$, vary across different ion channels and cell types. Steady-state mathematical analysis has shown that these differing time constants tune correlations in ion channel conductances. For two ion channels $i$ and $j$, steady-state analysis yields $\bar{g}_{i}/\bar{g}_{j}\approx\tau_{m_{j}}/\tau_{m_{i}}$. This indicates that homeostatic tuning rules maintain ratios of ionic conductances and their correlations while adjusting their values in response to intracellular calcium levels. In the ionic conductance space, these rules allow neurons to navigate along a one-dimensional line, where the slope of this line in two dimensions is determined by the ratios of mRNA transcription time constants. This phenomenon, known as homogeneous scaling, involves proportionally increasing or decreasing all ionic conductances by the same multiplicative factor.
On the other hand, sharp neuromodulation is implemented by transiently applying the algorithm of Fyon et al., (2024). This neuromodulation also produces changes in conductance values (Fig. 1 left), resulting in motion within the conductance plane in a direction different from homogeneous scaling (Fig. 1 right) to reach a new firing activity. However, neuromodulation only affects a subset of ion channels, not all (Marder,, 2012). Furthermore, the timescale of neuromodulation is much faster than that of homeostasis, lasting for a maximum of a few minutes (Marder et al.,, 2014). In addition, neuromodulation tunes correlations between neuromodulated ion channels, rather than preserving them (Fyon et al.,, 2024).
Coupling homeostasis and neuromodulation in this way is problematic because homeostasis preserves correlations and regulates calcium, whereas neuromodulation adjusts correlations to modulate neuronal function. In Fig. 1, sharp neuromodulation is applied to three different STG models with distinct initial ionic conductances. Initially, all neurons exhibited tonic firing due to homeostatic regulation, with RNA transcription time constants selected to produce tonic spiking neurons. Neuromodulation was then applied instantaneously, leading to a transition from tonic spiking to bursting in all cases (Fig. 1). These transitions drastically increase intracellular calcium concentration. From the perspective of homeostatic tuning rules, this constitutes a disturbance, as it alters the conductance ratios between $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$. Although this disturbance is necessary for adaptation and functional modulation, the outcome of homeostasis in response to this biological disturbance is unpredictable. Indeed, as the homeostatic controller was initially designed to produce tonic spiking neurons, it tends to preserve the predefined correlations that produce such behavior. However, neuromodulation changes the firing activity, leading to a mismatch between actual conductance values and the encrypted conductance ratios.
In Fig. 1A, homeostatic compensation results in pathological behavior, characterized by irregular bursting with a hyperpolarized plateau, failing to achieve either bursting or tonic spiking. In Fig. 1B, the modulated behavior (bursting) is maintained after homeostatic compensation. In Fig. 1C, homeostatic compensation leads to unphysiological values of $\bar{g}_{\mathrm{A}}$, tending towards negative values. Even when $\bar{g}_{\mathrm{A}}$ is constrained to be non-negative, the resulting electrical behavior is pathological, failing to exhibit either tonic spiking or bursting. This toy experiment highlights that homeostasis and sharp neuromodulation can interfere pathologically.
Such behavior was predicted in O’Leary et al., (2014), where the authors demonstrate that their homeostasis model can lead to various outcomes in response to disturbances, such as channel deletions. Initially, the deletion itself may directly affect neuronal function, either preserving it or causing its loss. Subsequently, homeostatic compensation can either preserve or restore function, but it may also lead to pathological behaviors, such as functional deletion. This result suggests that any disturbance applied to the conductance ratios—even physiological ones such as neuromodulation—can lead to problematic behaviors.
Figure 1: Current model of homeostasis can lead to pathological excitability in the presence of sharp neuromodulation. A. Time evolution of all conductances in the STG model, shown on a logarithmic scale, when homeostasis is applied with sharp neuromodulation (instantaneous changes in $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$ within the red region) (left). Corresponding trajectory in the modulated conductance space (right). The white circle represents the initial condition (tonic spiking), the red diamond marks the neuromodulated state (bursting) immediately after the sharp modulation, and the green triangle denotes the final state after homeostasis has converged. In this case, homeostatic compensation is pathological, resulting in irregular bursting. B. Same as in panel A, but with different degenerate initial conditions. Here, homeostatic compensation preserves the neuromodulated function. C. Same as in panel A, with another set of degenerate initial conditions. In this case, homeostatic compensation results in negative values of $\bar{g}_{\mathrm{A}}$ (saturated at 0), leading to pathological behavior characterized by one-spike bursting.
Controlled neuromodulation and homeostasis naturally lead to robust neuronal function
In biological systems, homeostasis and neuromodulation collaborate to ensure robust and modular neural function. These mechanisms operate on different timescales: neuromodulation acts within seconds or minutes (Werner and Mitterauer,, 2014), while homeostatic compensation may take hours or even days (Marder and Prinz,, 2002). However, in computational models, the goals of neuromodulation and homeostatic controllers often seem discordant. Neuromodulation typically adjusts conductance ratios, while homeostasis seeks to maintain these ratios and guide the neuron through homogeneous scaling. Consequently, applying sharp neuromodulation to any homeostatic conductance-based model can lead to instability, as neuromodulation disrupts the references and control actions of the homeostasis controller (Fig. 2A).
To model the actions of both mechanisms within a single cell, we include a more advanced neuromodulation model, referred to as the neuromodulation controller, as developed in Fyon et al., (2023). This controller can be coupled with any conductance-based model, allowing modulation of its neuronal function, such as its firing pattern, to mimic biological neuromodulation. We call this type of neuromodulation controlled neuromodulation, in contrast to the sharp neuromodulation discussed above. The neuromodulation controller adjusts target channel conductances, providing continuous and consistent modulation rather than instantaneous, rigid changes.
Fig. 2B illustrates the results of transitions from tonic spiking to bursting when applied to a degenerate population of STG neuron models. Initially, all neurons exhibit tonic firing mode, and homeostasis adjusts the conductances to achieve a predefined calcium level along a first tonic spiking homogeneous scaling line. Midway through the simulation, controlled neuromodulation is applied to switch all neurons to bursting mode, which now occurs through a controlled adjustment of $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$. Consequently, intracellular calcium levels rise sharply, and conductance ratios involving one or both neuromodulated ionic conductances are adjusted. Finally, with the reshaped RNA time constants, the homeostasis controller adjusts all conductances to restore an appropriate calcium level, while controlled neuromodulation preserves neuromodulated conductance ratios along a second bursting homogeneous scaling line. Ultimately, the neuronal population reaches a new state where the neuromodulated function is preserved and intracellular calcium levels are back to their target value.
Simulations were also performed on a midbrain dopaminergic neuron model (adapted from Qian et al., (2014)) to demonstrate the generality of our results (Fig. 2C). As with the previous simulations, a transition from tonic spiking (pacemaking) to strong bursting was induced midway through by modulating the N-type and L-type calcium conductances ($\bar{g}_{\mathrm{CaN}}$ and $\bar{g}_{\mathrm{CaL}}$) using controlled neuromodulation. Homeostatic compensation was continuously applied to maintain proper calcium levels, resulting in a reliable and robust outcome despite initial degeneracy within the population.
Figure 2: Controlled neuromodulation and homeostasis naturally lead to robust neuronal function. A. Schematic representation of neuromodulation (left) and homeostasis (right) cascades. Neuromodulators bind to G-protein receptors, triggering complex signaling pathways that selectively affect subsets of ion channels. In contrast, homeostasis senses intracellular calcium levels and regulates them by adjusting all conductances uniformly. B. Time evolution of all conductances in the STG model, displayed on a logarithmic scale, during homeostasis with controlled neuromodulation (controlled changes in $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$ starting at the dashed blue line) for a degenerate population of 200 models (left). The corresponding mean intracellular calcium concentration over time (right) shows the target value (red dash-dotted line) is maintained both before and after controlled neuromodulation. C. Same as panel B, but for a DA model where controlled neuromodulation affects $\bar{g}_{\mathrm{CaN}}$ and $\bar{g}_{\mathrm{CaL}}$.
Why do controlled neuromodulation and homeostasis lead to robust neural function?
Fig. 3A shows the trajectories of the simulations from Fig. 2B in the plane of modulated conductances, $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$, during tonic spiking homeostatic compensation for both sharp and controlled neuromodulation. Initially, the modulated conductances of every neuron in the population grow along the homogeneous scaling direction in tonic spiking mode, as the calcium levels are below the threshold value. This homogeneous scaling corresponds to a line intersecting the origin, with variability due to the degeneracy in conductance ratios within the population. During neuromodulation, all conductances are adjusted to tune their ratios and achieve bursting (Fig. 3B). Importantly, all neuromodulatory actions occur along the same direction, but with varying magnitudes depending on the state of the neuron (Fyon et al.,, 2024). As a result, calcium levels increase drastically for each neuron in the population. Once the neuromodulatory action ends, the outcome depends on the type of neuromodulation applied (Fig. 3C).
In the case of sharp neuromodulation, during the homeostatic action in the bursting phase, the homeostatic controller acts alone and tends to maintain the same conductance ratios as in tonic spiking (Fig. 3C left). This leads to a reduction in calcium levels, but, as seen in Fig. 1, its effects are unreliable and undesirable. Specifically, for most neurons in the population, the homeostatic controller reduces $\bar{g}_{\mathrm{A}}$ towards negative values (saturated at 0 to avoid model instability) to bring $\bar{g}_{\mathrm{CaS}}$ into an appropriate range. This regulation causes the neuron to move away from the robust bursting behavior, resulting in heterogeneous changes in the activity of the population.
In the case of controlled neuromodulation, the homeostatic and neuromodulation controllers work together to maintain the new conductance ratios corresponding to bursting (Fig. 3C right). This results in a new homogeneous scaling following these updated conductance ratios in bursting, which translates into trajectories that move toward the origin of the axes, as would be expected in tonic spiking. Homogeneous scaling during bursting differs across the population due to initial degeneracy. The difference between the two homogeneous scalings (tonic spiking and bursting) is driven by the neuromodulatory action, ensuring that the neuromodulated behavior (bursting) is not lost.
Figure 3: Why do controlled neuromodulation and homeostasis lead to robust neural function? A. Trajectories of the STG population from Fig. 2 in the modulated conductance space during homogeneous scaling in tonic spiking only. As expected from homeostasis, all neurons move in the same direction. White circles represent the initial conditions at the start of the step, and the green triangle marks the final conditions. B. Same as panel A, but during the brief period when neuromodulation is active. All trajectories are parallel to one another. C. Same as panels A and B, but during the homeostasis phase following neuromodulation. The outcome depends on the type of neuromodulation: sharp (left) or controlled (right). Sharp neuromodulation results in non-robust outcomes, whereas controlled neuromodulation preserves robust neural function.
In the first section, when the homeostasis controller was combined with sharp neuromodulation, unreliable outcomes were observed. To explain this behavior, we use a schematic representation of the modulated conductance plane (Fig. 4A). In this plane, different values of $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$ can lead to similar mean intracellular calcium concentration. These values can be highlighted using calcium isoclines (green dashed lines in Fig. 4A). Likewise, different values of $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$ can lead to similar firing activity. These values can be highlighted using activity isoclines (blue dashed lines in Fig. 4A). Calcium isoclines are parallel one to each other, whereas activity isoclines pass through the origin and rotate around this point. This behavior is explained by homeostatic tuning rules and homogeneous scaling, as different conductance ratios lead to different firing patterns. Initially, the neuron operates in the target spiking mode at the target calcium level, corresponding to the intersection of the spiking and the target calcium isoclines. When sharply neuromodulating the neuron, the modulated conductances change, and the neuron shifts to another state, now reaching a new target strong bursting isocline, which intersects a higher calcium isocline. Homeostasis then kicks in to reduce the calcium level to its target value. However, in the homeostatic controller, the “native” tonic spiking direction is encrypted, leading to changes in conductances along the spiking isocline. As a result, the neuron moves away from the bursting isocline, potentially disrupting the new function. This explains why sharp neuromodulation is incompatible with the homeostasis controller.
To harmonize the actions of both controllers and achieve robust functional modulation, we leverage the significant difference in their timescales. Neuromodulation controls the modulated ionic conductances, while homeostatic tuning rules regulate all ionic conductances over a slower timescale to maintain an appropriate calcium level. In the modulated conductance plane (Fig. 4B), the initial and neuromodulated states are the same as in Fig. 4A. However, in the case of controlled neuromodulation, both homeostatic and neuromodulatory controllers are active. Specifically, when homeostasis reduces the calcium level and shifts the neuron away from the bursting isocline, the neuromodulation action counteracts this perturbation in function by bringing the neuron back to the bursting isocline, preserving its function. The neuromodulatory controller operates several orders of magnitude faster than the homeostatic controller. As a result, fast corrections from the former render perturbations from the latter infinitesimal, leading to changes in the conductance ratio along the target activity isocline (strong bursting in this example). Ultimately, the neuron stabilizes at the intersection of the target calcium and target activity isoclines, where both controllers meet their objective.
Figure 4: Schematic explanation of why controlled neuromodulation integrates better with homeostasis compared to sharp neuromodulation. A. Schematic of the modulated conductance space for homeostasis combined with sharp neuromodulation. Initially (white circles), the neuron spikes at the target calcium level, located at the intersection of the firing pattern isocline (dashed blue lines) and the calcium level isocline (dashed green lines). Following sharp neuromodulation to a new firing pattern (red diamond), calcium levels sharply increase. Homeostasis reduces calcium (solid red arrow) by moving along the same direction as before neuromodulation (dashed red arrow), causing the neuron to deviate from the neuromodulated firing pattern isocline. B. Same as panel A, but with controlled neuromodulation. The initial steps are identical, but after neuromodulation, as homeostasis acts to reduce calcium, controlled neuromodulation simultaneously adjusts to keep the neuron on the neuromodulated firing pattern isocline (sawtooth red arrow). Because controlled neuromodulation operates on a much faster timescale than homeostasis, the neuron remains on the neuromodulated firing pattern isocline throughout.
Controlled neuromodulation and homeostasis ensure the preservation of function under physiologically recoverable disturbances
By employing this tandem of controllers, neuronal function becomes robust to channel blockade, provided that the channel deletion is compensable. While homeostasis compensation alone might lead to unreliable recovery from channel blockade, incorporating controlled neuromodulation somehow stabilizes the neuron response to such blockade, even in the case of highly degenerate neuronal populations. Furthermore, if the channel blockade is physiologically recoverable—meaning the function of the blocked channel can be restored using other channels—there exists an optimal combination of neuronal feedback gains to preserve that function. In essence, if one channel is blocked, a specific concentration of particular neuromodulators can preserve the function.
For example, in Fig.5, various channel blockades are applied to a degenerate population of STG conductance-based models equipped with the combined homeostasis/neuromodulation tandem of controllers. The target behavior of the population is regular bursting. In Fig.5A, H-type channels are blocked. Here, the function is preserved even immediately after the blockade, without requiring compensation. Subsequently, the combined regulation of homeostasis and controlled neuromodulation maintains this function by restoring the neuron to an appropriate calcium level on a population-wide scale. In Fig. 5B and C, T-type calcium and calcium-controlled potassium channels are blocked, respectively. In these cases, the blockade results mostly in the loss of bursting function. However, by allowing the homeostasis/neuromodulation pair of controllers to compensate, the function is largely recovered, albeit with irregular bursting patterns in some neurons. Additionally, intracellular calcium levels stabilize at their target values. This preservation of functions is feasible because the actions of H-type, T-type calcium, and calcium-controlled potassium channels are degenerate and can be compensated for using other channels present within the cell. The STG neuron model incorporates numerous channels operating on slow and ultraslow timescales, enabling the compensation for lost channels in these timescales. It is worth noting that only the loss of the calcium-controlled potassium channel has been compensated using different neuronal feedback gain values, i.e., different concentrations of various neuromodulators, as in control conditions.
However, the blockade of fast transient sodium channels is irrecoverable (Fig. 5D), as these channels are crucial for initiating spikes on a fast timescale. Once the blockade is applied, compensation cannot restore spikes because the action of sodium channels is non-degenerate and occurs on the fast timescale of the neuron. Similarly, achieving the appropriate intracellular calcium level becomes unattainable. A comparable outcome would result from the blockade of delayed rectified potassium channels, which are responsible for the downstroke of the spikes.
Figure 5: Controlled neuromodulation and homeostasis ensure the preservation of function under physiologically recoverable disturbances. A. Time evolution of all conductances in the STG model, displayed on a logarithmic scale, during homeostasis with controlled neuromodulation (controlled changes in $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$ starting at the dashed blue line) for a degenerate population of 200 models with H-type ion channel blockade (left). The corresponding distribution of behaviors before the blockade, immediately after the blockade, and following compensation (right) demonstrates the effect of the blockade. In this case, no neurons lose bursting due to the channel blockade or compensation. B. Same as panel A, but with a CaT channel blockade. The blockade causes half of the neurons to lose bursting; however, all neurons recover bursting with compensation. C. Same as panel A, but with a KCa channel blockade. The blockade causes more than half of the neurons to lose bursting, but nearly all recover bursting with compensation. D. Same as panel A, but with a Na channel blockade. The blockade causes all neurons to lose bursting, and compensation does not restore bursting due to the non-degenerate essential role of Na channels.
Central pattern generator networks can be robustly modulated to orchestrate their rhythmic output
The application of this tandem of controllers extends even to networks, particularly central pattern generators (CPGs). CPGs are self-organized neuronal circuits mainly composed of bursting neurons connected primarily through inhibitory connections, capable of generating rhythmic output without external input (Yuste et al.,, 2005; Guertin,, 2009). One extensively studied circuit is responsible for the pyloric and gastric mill rhythms in crustaceans, observed in the stomatogastric ganglion. This network exhibits two rhythms at different speeds: the fast pyloric rhythm and the slow gastric mill rhythm. These rhythmic outputs are vital for coordinating muscle contractions in the stomach for digestion, among other functions. A minimalist network model of such a bi-rhythmic circuit is investigated in Fig. 6 (Gutierrez et al.,, 2013; Drion et al.,, 2019). It comprises two half-center oscillators operating at different speeds, connected via a central neuron. Half-center oscillators consist of two bursting neurons connected in mutual inhibition, enabling them to fire anti-phase bursts. In the network scheme, the purple and blue neurons constitute the fast half-center oscillator (representing the pyloric rhythm), while the green and yellow neurons form the slow half-center oscillator (representing the gastric mill rhythm). The red neuron serves as the central one, also in bursting mode. It has been demonstrated that the gastric mill rhythm requires a neuromodulatory input to be activated (Selverston et al.,, 2009).
In this study, the gastric mill rhythm can be reliably induced by the pair of controllers, but not with homeostasis combined with sharp neuromodulation. Under control conditions (Fig. 6A), i.e., without neuromodulatory input to the gastric mill rhythm, only the pyloric rhythm is active and the central neuron consistently bursts. In particular, variability in conductances is observed in this network (Fig. 6A right). When neurons are equipped with homeostasis only, activation of sharp neuromodulation for the gastric mill rhythm leads to rapid loss of function, due to the unreliability of this combination (Fig. 6B middle). However, when neurons undergo controlled neuromodulation, the gastric mill rhythm is successfully induced and sustained (Fig. 6B right). Despite the degeneracy of neurons, characterized by significant variability in conductance ratios, the network achieves rhythmic output with appropriate intracellular calcium levels. This toy experiment demonstrates that such a tandem of controllers can be applied in computational models to modulate and maintain not only single cell activity but also network-level activity.
Figure 6: Central pattern generator networks can be robustly modulated to orchestrate their rhythmic output. A. Control condition of the simplified pyloric/gastric mill rhythm network (left), inspired from Gutierrez et al., (2013). Without neuromodulation, the fast rhythm is active (top two traces, middle), while the slow rhythm is inactive (bottom two traces, middle). In this case, all neurons in the network exhibit degenerate conductances (right). B. Schematic representation of the activated slow rhythm network (left), modulated by either sharp neuromodulation (middle) or controlled neuromodulation (right). With sharp neuromodulation, the slow rhythm fails due to its unreliability, whereas controlled neuromodulation sustains the slow rhythm.
DISCUSSION
The efficient and robust integration of homeostasis and neuromodulation in neurons remains an open question with several proposed hypotheses (Marder et al.,, 2014; Marder and Bucher,, 2005; Cunha,, 2001). The improper integration of these mechanisms—for example, combining sharp neuromodulation with a homeostatic controller—can lead to negative interactions. Such interactions may result in unreliable behaviors or numerical instabilities.
In contrast, pairing a homeostatic controller with controlled neuromodulation ensures reliable neural function modulation while maintaining appropriate calcium levels. This approach draws inspiration from biology, where neuromodulation is known to be activity-dependent, as observed experimentally (Kramer and Levitan,, 1990; Raymond et al.,, 1992). Specifically, the neuromodulation controller models the role of G-protein-coupled receptors and second messengers, embedding feedback in the metabolic neuromodulation cascade (Marcus and Bruchas,, 2023; Scheler,, 2004). This reliable interaction can be extended to neural circuits, such as the gastric mill circuit in crabs, to produce robust network-level functionality (Marder et al.,, 2014).
However, the success of this interaction critically depends on a point in the conductance space where both controllers converge. This point represents a neural activity pattern that supports neuromodulated firing while maintaining homeostatic calcium levels. As shown in Fig. 4, this intersection of the strong bursting isocline (the target of controlled neuromodulation) and the target calcium level (the goal of homeostatic control) represents the convergence point. When such a point is absent—e.g., during sodium channel blockade (Fig. 5)—neither neuromodulated activity nor calcium homeostasis is sustained. This may lead to pathological compensations, where one controller prioritizes calcium regulation at the cost of abandoning the desired firing pattern (Starmer et al.,, 1991; Grant et al.,, 1984; Anger et al.,, 2001). Furthermore, discontinuities in the neuromodulated firing pattern may lead to transient pathological behaviors.
Maximizing convergence between these control mechanisms requires increasing the likelihood of both an intersection and a continuous path in conductance space. We hypothesize that this can be achieved by maximizing neuronal degeneracy (Whitacre,, 2010; Rathour and Narayanan,, 2019; Whitacre and Bender,, 2010; Prinz,, 2017; Goaillard and Marder,, 2021). Higher degeneracy increases the likelihood that the functional and dynamical role of one ion channel can be compensated by others, thereby preserving both the path and the intersection. This underscores the essential role of degeneracy in enabling robust neural functions that can be modulated.
Finally, this study underscores the potential risks associated with pharmacological ion channel blockade (X Cubeddu,, 2016; DeWitt and Waksman,, 2004; Imbrici et al.,, 2018). If neuronal degeneracy is insufficient, such interventions may disrupt neuron ability to maintain robust neuromodulation. Instead, we propose targeting elements within the neuromodulation cascade—such as second messengers or other components of the control loop—as more reliable pharmacological strategies. Such interventions would preserve the dynamics of controlled neuromodulation, providing a more robust alternative to direct ion channel blockade.
MATERIALS AND METHODS
Programming language
The Julia programming language was used in this work Bezanson et al., (2017). Numerical integration was realized using DifferentialEquations.jl.
Conductance-based models
For all experiments, single-compartment conductance-based models were employed. These models articulate an ordinary differential equation for the membrane voltage $V$, where $N$ ion channels are characterized as nonlinear dynamic conductances, and the phospholipid bilayer is represented as a passive resistor-capacitance circuit. Mathematically, the voltage-current relationship of any conductance-based neuron model is expressed as follows:
$I_{C}$ $=C\frac{\mathrm{d}V}{\mathrm{d}t}+g_{\mathrm{leak}}(V-E_{\mathrm{leak}})=-I_{\mathrm{int}}+I_{\mathrm{ext}},$
$=-\sum_{\mathrm{ion}\in\mathcal{I}}g_{\mathrm{ion}}(V,t)(V-E_{\mathrm{ion}})+I_{\mathrm{ext}}.$
Here, $C$ represents the membrane capacitance, $g_{\mathrm{ion}}$ denotes the considered ion channel conductance and is non-negative, gated between 0 (all channels closed) and $\bar{g}_{\mathrm{ion}}$ (all channels open), $E_{\mathrm{ion}}$ and $E_{\mathrm{leak}}$ are the channel reversal potentials, $\mathcal{I}$ is the index set of intrinsic ionic currents considered in the model, and $I_{\mathrm{ext}}$ is the current externally applied in vitro, or the combination of synaptic currents. Each ion channel conductance is nonlinear and dynamic, represented by $g_{\mathrm{ion}}(V,t)=\bar{g}_{\mathrm{ion}}m_{\mathrm{ion}}^{a}(V,t)h_{\mathrm{ion}}^{b}(V,t)$, where $m_{\mathrm{ion}}$ and $h_{\mathrm{ion}}$ are variables gated between 0 and 1, modeling the opening and closing gates of ion channels, respectively. Throughout this study, both the isolated crab STG neuron model Liu et al., (1998) and the adapted DA neuron model Qian et al., (2014) (where SK channels had been blocked to enable bursting) were employed.
The STG model consists of seven ion channels that operate on various time scales: fast sodium channels ($\bar{g}_{\mathrm{Na}}$); delayed-rectifier potassium channels ($\bar{g}_{\mathrm{Kd}}$); T-type calcium channels ($\bar{g}_{\mathrm{CaT}}$); A-type potassium channels ($\bar{g}_{\mathrm{A}}$); slow calcium channels ($\bar{g}_{\mathrm{CaS}}$); calcium controlled potassium channels ($\bar{g}_{\mathrm{KCa}}$); and H channels ($\bar{g}_{\mathrm{H}}$).
The DA model consists of six ion channels that operate on various time scales: fast sodium channels ($\bar{g}_{\mathrm{Na}}$); delayed-rectifier potassium channels ($\bar{g}_{\mathrm{Kd}}$); L-type calcium channels ($\bar{g}_{\mathrm{CaL}}$); N-type calcium channels ($\bar{g}_{\mathrm{CaN}}$); ERG channels ($\bar{g}_{\mathrm{ERG}}$); and NMDA channels ($\bar{g}_{\mathrm{NMDA}}$). Note that, owing to the multicellular nature of NMDA channels, they were excluded from this study but were still used for simulations with baseline values.
The homeostatic controller
The homeostatic controller adjusts all conductances using an integration rule described in O’Leary et al., (2014). Let $\bar{g}_{i}$ denote the conductance of ion channel $i\in[1,N]$, and $m_{i}$ the corresponding mRNA concentration. The dynamics of the homeostatic controller are governed by the following equations:
$\tau_{i}\dot{m}_{i}$ $=[Ca^{+2}]_{\mathrm{target}}-[Ca^{+2}],$
$\tau_{g}\dot{\bar{g}}_{i}$ $=m_{i}-\bar{g}_{i},$
where $[Ca^{+2}]$ and $[Ca^{+2}]_{\mathrm{target}}$ represent the intracellular calcium concentration and its target value, respectively. In essence, the calcium concentration error is integrated into the mRNA levels, which subsequently modulate the corresponding channel conductances. This regulation aims to bring the calcium concentration to its target value, as higher channel conductances (particularly calcium-conducting ones) facilitate greater calcium influx. Starting from any initial condition, this controller drives the cell along a line in conductance space known as the homogeneous scaling line. This line passes through both the initial condition and the origin of the conductance space axes. Steady-state analysis reveals that the conductance ratios remain invariant, such that $\frac{\bar{g}_{i}}{\bar{g}_{j}}=\frac{\tau_{j}}{\tau_{i}}$. This invariance ensures that the relative scaling of conductances is preserved during homeostatic compensation. Consequently, the cell’s conductance properties develop along the homogeneous scaling beam defined by these ratios.
The neuromodulation controller
The neuromodulation controller adjusts a subset of $n$ conductances, denoted as $\bar{g}_{\mathrm{mod}}\in\mathbb{R}^{n}$ with $n<N$, using the algorithm introduced in Fyon et al., (2023). This controller employs the concept of Dynamic Input Conductances (DICs) to determine target values for $\bar{g}_{\mathrm{mod}}$ based on the input neuromodulator concentration, represented as $\bar{g}_{0}\left([\mathrm{nmod}]\right)\in\mathbb{R}^{n}$. The error between the target and the current values, $e_{\mathrm{mod}}$, drives a Proportional-Integral (PI) controller, which updates $\bar{g}_{\mathrm{mod}}$ to track the desired reference. The controller dynamics are described by the following equations:
$e_{\mathrm{mod}}$ $=\bar{g}_{0}\left([\text{nmod}]\right)-\bar{g}_{\mathrm{mod}},$
$\dot{\bar{g}}_{\mathrm{mod}}$ $=f\left(K_{p}\cdot e_{\mathrm{mod}}+K_{i}\cdot\int e_{\mathrm{mod}}\mathrm{d}t\right),$
where $K_{p}$ and $K_{i}$ are the proportional and integral gains of the PI controller, respectively. In summary, the controller adjusts the modulated conductances $\bar{g}_{\mathrm{mod}}$ in response to the local concentration of neuromodulator, producing new reference values for these conductances to fine-tune the cell’s firing pattern—effectively neuromodulating it. This neuromodulator controller produces changes in conductance ratios that are normally preserved within the homeostatic controller. The new reference values depend on the neuromodulator concentration, the type of neuron, and the cell’s current state, characterized by the full set of conductances $\bar{g}_{\mathrm{ion}}\in\mathbb{R}^{N}$. The computation of these references is achieved through DICs, which encapsulate this complexity.
DICs consist of three voltage-dependent conductances that separate according to timescales: one fast, one slow, and one ultraslow, denoted as $g_{\mathrm{f}}(V)$, $g_{\mathrm{s}}(V)$, and $g_{\mathrm{u}}(V)$, which can be computed as linear functions of the maximal conductance vector $\bar{g}_{\mathrm{ion}}\in\mathbb{R}^{N}$ of an $N$-channel conductance-based model at each voltage level $V$:
$[g_{\mathrm{f}}(V);g_{\mathrm{s}}(V);g_{\mathrm{u}}(V)]=f_{\mathrm{DIC}}(V)=S(V)\cdot\bar{g}_{\mathrm{ion}}\,,$
where $S(V)\in\mathbb{R}^{3\times N}$ is a sensitivity matrix that can be built by: $S_{i\,j}(V)=-\left(w_{i\,j}\cdot\frac{\partial\dot{V}}{\partial X_{j}}\frac{\partial X_{j,\infty}}{\partial V}\right)/g_{\mathrm{leak}}$, where $i$ denotes the timescale, $X_{j}$ are gating variables of the j-th channel of the considered model and $w_{i\,j}$ is a timescale-dependent weight which is computed as the logarithmic distance of the time constant of $X_{j}$ and the timescale $i$ (Drion et al., 2015a, ). While the complete curve of the DICs may be of interest, only its value at the threshold voltage $V_{\mathrm{th}}$ is used, as the values and signs of the DICs at $V_{\mathrm{th}}$ reliably determine the firing pattern Drion et al., 2015a . Thus, the following linear system $f_{\mathrm{DIC}}(V_{\mathrm{th}})=S(V_{\mathrm{th}})\cdot\bar{g}_{\mathrm{ion}}$ makes the link between ion channel conductances and neuronal activity.
COMPETING INTEREST
The authors declare no competing interests.
FUNDING
Arthur Fyon is a Research Fellow of the “Fonds de la Recherche Scientifique - FNRS”, supported by grant ASP-REN40024838. This work was supported by the Belgian Government through the Federal Public Service (FPS) Policy and Support (BOSA) grant NEMODEI (NEuroMOrphic Design of Embodied Intelligence).
AUTHOR CONTRIBUTIONS STATEMENT
A.F. and G.D. conceived and designed the research. A.F. and G.D. developed the model code. A.F. performed the simulations and analyzed the data. A.F. and G.D. interpreted the results. A.F. prepared figures and drafted the manuscript. A.F. and G.D. edited and revised the manuscript. A.F. and G.D. approved the final version of the manuscript.
DATA AVAILABILITY
All code and data can be found on the first author’s GitHub (https://github.com/arthur-fyon/BIOCONTROL_2025).